Question 46 of 75. Philip is a single taxpayer with modified AGI of $135,000 and an active participation rental real estate loss of $30,ooo. Assuming he has no other source of passive income, what is his adjusted special loss allowance? Oso o $5,000 o $7,500 O $25,000
Answer #1

solution :

given that

Philip is a single taxpayer with modified AGI of $135000 and an active participation rental real estate loss of $30000

also given that assuming

he has no other source of passive income ,

calculation of adjusted special loss allowance :

Phase out = [(135000-100000)/50000] x 25000 = 17500

Allowed loss = 25000-17500 = 7500

Explanation:

Misfortune from detached action is for the most part permitted to be set off against inactive movement salary, yet as a special case to this standard misfortune from rental land with dynamic support is permitted to be set off upto $25,000. There is pertinent eliminate to this $25,000 setoff if MAGI surpasses $100,000 and comes to $150,000. For example each $2 surpasses MAGI $100,000, $1 sum is decreased.

therefore

option c

$7500